# Granular-ball Guided Masking: Structure-aware Data Augmentation

### v1

@@ IV-C Empirical Complexity Analysis
Results in Table VII show Random Erasing is fastest and HAS slowest, with AutoAugment/GridMask moderate. GBGM averages 3.783.78ms and 2256.3 samples/sec. Table VIII further shows $\beta=0.601$, meaning that per-sample cost decreases as batch size grows. This observation is consistent with our theoretical analysis: although GBGM has linear complexity $O(HW)$ with respect to image resolution, its mask generation pipeline is highly parallelizable, allowing GPUs to amortize computation across larger batches. As a result, GBGM achieves favorable sub-linear scalability in practice, making it particularly suitable for large-scale parallel training.
TABLE VII: Runtime performance of augmentation methods (ms and samples/sec).
@@
Random Erasing [zhong2020random] 0.14 3.70 1.21 8348.3
AutoAugment [cubuk2019autoaugment] 0.82 20.51 6.65 1547.6
GridMask [chen2020gridmask] 0.47 13.59 4.52 2322.3
HAS [singh2018hide] 1.30 39.12 12.87 810.3
@@
TABLE VIII: Scalability factors of different augmentation methods.
@@
Random Erasing [zhong2020random] 0.950 ($R^{2}=0.997$) Linear
AutoAugment [cubuk2019autoaugment] 0.956 ($R^{2}=0.993$) Linear
GridMask [chen2020gridmask] 0.994 ($R^{2}=0.998$) Linear
HAS [singh2018hide] 0.991 ($R^{2}=1.000$) Linear
@@ V Conclusion
We present Granular-ball Guided Masking (GBGM), a structure-aware augmentation strategy that unifies granular-ball computing with hierarchical masking. GBGM adaptively preserves informative regions while suppressing redundancy, yielding coarse-to-fine masks for both CNNs and Vision Transformers. Experiments on CIFAR-10/100, ImageNet-100, and ImageNet-1K show consistent accuracy gains and favorable sub-linear scalability. Looking ahead, GBGM offers a promising direction for self-supervised pretraining and dense prediction tasks, where structure-aware masking supports spatial and temporal consistency.

### v2

@@ IV-D Empirical Complexity Analysis
Results in Table VIII show Random Erasing is fastest and HAS slowest, with AutoAugment/GridMask moderate. GBGM averages 3.783.78ms and 2256.3 samples/sec. Table IX further shows $\beta=0.601$, meaning that per-sample cost decreases as batch size grows. This observation is consistent with our theoretical analysis: although GBGM has linear complexity $O(HW)$ with respect to image resolution, its mask generation pipeline is highly parallelizable, allowing GPUs to amortize computation across larger batches. As a result, GBGM achieves favorable sub-linear scalability in practice, making it particularly suitable for large-scale parallel training.
TABLE VIII: Runtime performance of augmentation methods (ms and samples/sec).
@@
Random Erasing [34] 0.14 3.70 1.21 8348.3
AutoAugment [5] 0.82 20.51 6.65 1547.6
GridMask [3] 0.47 13.59 4.52 2322.3
HAS [24] 1.30 39.12 12.87 810.3
@@
TABLE IX: Scalability factors of different augmentation methods.
@@
Random Erasing [34] 0.950 ($R^{2}=0.997$) Linear
AutoAugment [5] 0.956 ($R^{2}=0.993$) Linear
GridMask [3] 0.994 ($R^{2}=0.998$) Linear
HAS [24] 0.991 ($R^{2}=1.000$) Linear
@@ V Conclusion
We presented Granular Ball Guided Masking (GBGM), a structure-aware augmentation strategy that integrates Granular Ball Computing with hierarchical masking. By organizing image regions into a coarse-to-fine structure, GBGM preserves semantically informative and structurally important content while suppressing redundant areas, producing discriminative masks that can be seamlessly applied to both CNNs and Vision Transformers. Extensive experiments on CIFAR-10/100, ImageNet-100, and ImageNet-1K demonstrate consistent gains in classification accuracy and favorable sub-linear scalability. Beyond recognition, GBGM also improves performance on image tampering detection under the ADCD-based setting, suggesting that structure-aware masking can benefit forensic robustness under distribution shifts. Future work will explore extending GBGM to self-supervised pretraining and dense prediction tasks, where structure-aware masking may further support spatial and temporal consistency.